# MessIRve: A Large-Scale Spanish Information Retrieval Dataset

### v1

@@ Abstract
Information retrieval (IR) is the task of finding relevant documents in response to a user query. Although Spanish is the second most spoken native language, current IR benchmarks lack Spanish data, hindering the development of information access tools for Spanish speakers. We introduce MessIRve, a large-scale Spanish IR dataset with around 730 thousand queries from Google’s autocomplete API and relevant documents sourced from Wikipedia. MessIRve’s queries reflect diverse Spanish-speaking regions, unlike other datasets that are translated from English or do not consider dialectal variations. The large size of the dataset allows it to cover a wide variety of topics, unlike smaller datasets. We provide a comprehensive description of the dataset, comparisons with existing datasets, and baseline evaluations of prominent IR models. Our contributions aim to advance Spanish IR research and improve information access for Spanish speakers.
@@
Francisco Valentini1, Viviana Cotik1,2, Damián Furman1, Ivan Bercovich3, Edgar Altszyler4, Juan Manuel Pérez1 1CONICET-UBA. Instituto de Ciencias de la Computación. Buenos Aires, Argentina 2UBA. FCEyN. Departamento de Computación. Buenos Aires, Argentina 3University of California, Santa Barbara. California, U.S.A. 4Quantit {fvalentini,jmperez,vcotik}@dc.uba.ar, damian.a.furman@gmail.com, ibercovich@ucsb.edu, edgar@bequantit.com
@@ 1 Introduction
Given a user query, Information retrieval (IR) is the process of locating relevant documents from a collection of data. IR systems have become increasingly important as they are the backbone for the currently widespread Retrieval Augmented Generation (RAG) systems, in which IR provides relevant passages that are subsequently fed into a Large Language Model (LLM). This approach has been shown to improve the quality of generated text, reducing the likelihood of factual errors or hallucinations by LLMs in several tasks (Cheng et al., 2023; Jiang et al., 2023; Cheng et al., 2024; Lin et al., 2024).
Natural Language Processing (NLP) research relies heavily on evaluation datasets, which establish a common ground for comparing retrieval algorithms and guide the development of new methods.
While a wealth of resources is available for English, other languages often rely on multilingual models and datasets. When addressing tasks in Spanish, the resources are notably scarce. Spanish is the the second most spoken language in the world by number of native speakers, with over 486 million speakers. It is also the official language of 20 countries and is widely spoken in the United States (Eberhard et al., 2024).
Despite its prevalence, at the time of writing this paper, the retrieval section of the popular MTEB benchmark (Muennighoff et al., 2023) includes datasets for English, Chinese, French, and Polish, but omits Spanish.111huggingface.co/spaces/mteb/leaderboard
The lack of good quality evaluation datasets is a major challenge for the development of IR systems for Spanish. Which systems are best for Spanish IR? This key question remains unanswered due to the lack of a comprehensive Spanish IR benchmark.
@@
    Introducing MessIRve, a new large-scale dataset in Spanish that accounts for the varieties of Spanish spoken accross the Spanish-speaking countries. [^2] MessIRve means works for me in Spanish (“me sirve”). The reference to Lionel Messi, player of the most popular sport in Spanish-speaking countries, football, also stresses the importance of using topics that are relevant to Spanish speakers.
@@
    Providing baseline evaluations of relevant existing models on the new dataset.
@@
    Making the dataset publicly available so that other researchers can use it to develop and evaluate IR systems for Spanish. [^3] The dataset is available at https://huggingface.co/datasets/spanish-ir/google_qrels under an appropriate license.
These contributions aim to spur further research in Information Retrieval for the Spanish language and to facilitate the development of efficient information access tools for Spanish speakers.
@@ 2 Related work
mMARCO (Bonifacio et al., 2021) 8 841 823 532 761 1.00 - - 6 979 1.07
SQAC (Gutiérrez-Fandiño et al., 2022) 6 247 14 934 1.01 1 861 1.00 1 908 1.00
MIRACL (Zhang et al., 2023) 10 373 953 2 162 9.96 648 9.94 1 515 9.95
PRES (Kamateri et al., 2019) 10 037 - - - - 167 7.72
Multi-EuP (Yang et al., 2023) 2 371 - - - - 633 3.75
MessIRve (full) 14 047 759 571 120 1.05 - - 160 099 1.02
@@
In this section, we review the Spanish IR datasets that are well-documented and publicly accessible. Relevant statistics of these datasets are presented in Table 1.
Following standard terminology, we refer to a dataset as a collection of queries, documents, and relevance judgments (Thakur et al., 2021). A query is a user’s need for information, a document is a passage that may contain the information the user seeks, and a relevance judgment indicates whether a document is relevant to a query. The collection of all documents is referred to as a corpus.
The multilingual MS MARCO (mMARCO, Bonifacio et al., 2021) is a multilingual version of the well-known MSMARCO dataset (Bajaj et al., 2016), translated into Spanish using both open-source machine translation and Google Translate. Even if not originally in Spanish, this translated dataset is a substantial resource for training IR models since it leverages the large set of questions sampled from Bing’s search query logs which make up the original dataset. For each question, annotators judged the relevance of around 10 passages retrieved by Bing’s IR system. The collection of documents is the union of all retrieved passages.
A non-translated dataset is the Spanish Question Answering Corpus (SQAC, Gutiérrez-Fandiño et al., 2022). Inspired by SQUAD (Rajpurkar et al., 2016), Spanish-speaking annotators were shown passages from the Spanish Wikipedia, Wikinews, and the AnCora corpus, and were asked to come up with up to five questions for each fragment, encouraging the use of synonyms to prevent lexical overlap. Even if it is not a true IR dataset because it lacks a collection of documents, the set of all relevant passages can be used as a corpus for IR.
MIRACL (Multilingual Information Retrieval Across a Continuum of Languages, Zhang et al., 2023) is a multilingual dataset that includes Spanish and 17 other languages, focusing on good-quality relevance judgments by native speakers. Annotators were given a random paragraph from a Wikipedia article and were asked to formulate related questions, which may or may not be answerable by the passage. For each question, other annotators assessed the relevance of 10 candidate passages retrieved by three different IR models. Although MIRACL contains a relatively small number of queries, it benefits from having more annotations per query than other datasets, and from the inclusion of negative relevance judgments as well.
In addition to these multi-domain datasets, we identified two datasets focusing on specific domains. The Spanish Passage Retrieval dataset (PRES, Kamateri et al., 2019) contains documents representing articles from reliable Spanish health websites. The dataset is made up of health-related queries and employs a pooling method for relevance annotations, whereby human annotators assess the relevance of the pooled, top-ranked documents retrieved by different IR systems. Each document is a health-related article.
Finally, the Multilingual European Parliament Dataset (Multi-EuP, Yang et al., 2023) leverages official debate topics from the European Parliament. Each topic is a query, and the documents are speeches from Spanish parliamentarians that took part in the debate. The relevance judgments indicate whether a speech was part of the debate on the given topic or not.
@@ 3.1 Data collection
MessIRve is built by using questions from Google’s autocomplete API (www.google.com/complete) as queries, and answers from Google Search “featured snippets” that link to Wikipedia as relevant documents. This data collection strategy is inspired by the GooAQ dataset (Khashabi et al., 2021).
@@
Starting with 120 predefined prefixes such as “qué” (what), “cómo” (how), “dónde” (where), “quién” (who), etc., we used the Google autocomplete API to obtain popular queries in Spanish starting with these prefixes (refer to Appendix A, Table 3 for the list of prefixes). We then iteratively expanded the set of prefixes based on the returned results until we reached a predefined number of results. The API does not always return queries that begin strictly with the provided prefixes, resulting in a dataset where not all queries begin with the initial prefixes.
@@
To find relevant passages for each query, we obtained “featured snippets” from Google Search that sourced from Wikipedia. These snippets are text excerpts displayed at the top of the search results page in response to a query, deemed by Google to be particularly relevant to the user’s information need. In 2018 these snippets were reported to have a 97.4% accuracy rate in terms of providing relevant information. [^4] https://blog.google/products/search/reintroduction-googles-featured-snippets/
@@
We built a Spanish Wikipedia corpus with each paragraph corresponding to a single document. For this we used WikiExtractor (Attardi, 2015) to process the Spanish Wikipedia dump of 2024-04-01 and matched the extracted Wikipedia passages to the valid (query, document) pairs obtained from Google Search featured snippets.
@@ 3.2 Data splits
Country Code #Q #Q #Unique Rel. Docs #Unique Rel. Articles Avg Q Length Avg Rel. Doc Length
- full 571 120 160 099 62 739 48 327 5.8 80.3
- none 358 014 99 385 43 490 34 745 5.8 81.1
Argentina ar 22 261 5 780 3 985 3 724 5.5 80.1
Bolivia bo 25 015 4 707 3 181 2 892 5.4 80.4
Chile cl 22 330 5 564 3 781 3 526 5.4 79.5
Colombia co 25 589 5 992 3 978 3 684 5.7 80.5
Costa Rica cr 23 605 5 747 4 111 3 839 5.5 80.5
Cuba cu 21 805 5 053 3 494 3 265 5.5 81.3
Dominican Rep. do 27 470 5 719 3 816 3 486 5.7 81.4
Ecuador ec 27 289 6 384 4 346 3 944 5.9 83.5
Spain es 23 115 7 509 5 205 4 913 5.6 80.1
Guatemala gt 23 311 4 290 2 959 2 726 5.6 80.5
Honduras hn 26 736 5 690 3 939 3 675 5.6 80.5
Mexico mx 31 957 8 400 5 618 5 198 5.6 79.7
Nicaragua ni 28 064 5 902 4 008 3 730 5.5 79.1
Panama pa 25 760 5 801 4 121 3 834 5.5 80.8
Peru pe 25 531 5 804 3 948 3 665 5.5 80.6
Puerto Rico pr 26 672 6 280 4 388 4 098 5.5 80.6
Paraguay py 24 717 5 474 3 867 3 574 5.5 78.6
El Salvador sv 25 920 5 821 4 078 3 785 5.6 79.6
United States us 23 197 4 535 2 942 2 784 5.6 81.1
Uruguay uy 20 856 5 571 4 062 3 816 5.4 77.4
Venezuela ve 26 730 6 126 4 091 3 715 5.7 82.8
Table 2: The MessIRve dataset. #: number of, Q: queries, Rel.: relevant, Doc.: document. Queries come from Google’s autocomplete API and documents are Wikipedia paragraphs. Therefore, a paragraph may be relevant to multiple queries. The full set includes queries that are not specific to any country (code=none) plus the set of unique queries from all countries. Each query has one relevant document in each country, while in the full set, a query may have multiple relevant documents. The statistics for the training set are nearly identical to those of the test set and are thus omitted.
The dataset is partitioned into training and test queries in such a way that the Wikipedia article to which the paragraph belongs is present in only one of the splits. The partitioning was done by country, with about 20% of the articles assigned to the test set. This split by article aims to reduce the overlap in topics between the training and test sets.
Queries not specific to any country (label none) were combined with the set of unique queries from all countries to form a full set, which is also split into training and test sets. Unlike the country-specific sets, in the full set some queries can have multiple relevant documents because the same query may return different featured snippets in different country domains. Statistical information about the dataset splits is provided in Table 2.
@@ 3.3 Quality assessment
We evaluated the quality of the dataset using a method inspired by Rybak (2023), sampling query-document pairs from the full test set and manually rating pairs on three binary criteria:
@@
    Document relevance: Does the document help answer the query? Even if the passage does not provide an explicit and thorough answer, it can be considered relevant if it contains information that, when combined with other sources, could help answer the query.
Ratings were performed by four unpaid native Spanish speakers from Argentina. Each rater evaluated 200 query-document pairs, with 50 pairs being common across all raters. This resulted in a total of 650 unique rated pairs.
93.5% of the queries were rated as correct, 90.1% were rated as unambiguous, and 90.2% of the documents were considered relevant to their queries, using the majority vote for overlapping queries. These results show that the dataset’s quality is acceptable, as the queries are generally clear and the documents annotated as relevant are most likely to contain information that helps answer the query.
Because these three criteria are intrinsically ambiguous, i.e. they depend on each rater’s interpretation, we calculated the agreement amongst raters by computing the percentage of the 50 common queries on which all annotators agree. The percentages are: 78% for query correctness, 74% for query unambiguity, and 78% for document relevance. These results support the reliability of the quality assessment.
@@ 4 Comparison with existing datasets
A key feature of MessIRve is the detailed documentation of its data collection process. Existing open-domain datasets, such as MIRACL and SQAC, even if they are relatively well-documented, lack detailed descriptions of how their Wikipedia passages were selected before annotation.
MessIRve is also the only dataset that aims at explicitly accounting for the diverse varieties of Spanish spoken in different countries. This contrasts with all other datasets, which either do not consider dialectal variations or do not provide clear information about their inclusion.
To validate the dataset’s dialectal diversity, we listed the most distinctive words from each country’s queries according to a frequency-based log odds ratio score (see Table 4 in Appendix A). Each country’s queries include terms that are specifically relevant to that country, such as “green” for the United States (referring to the Green Card), “checar” for Mexico (meaning “to check”), “frailejones” for Colombia (a type of plant), “euskera” for Spain (the Basque language), and “CUIT” for Argentina (a tax identification number), when considering the five countries with the most speakers.
@@
Moreover, MessIRve is substantially larger than other datasets (see Table 1). The mMARCO dataset, while large, involves translations from English, thus introducing potential translation errors and artifacts which may affect the quality of the dataset. Moreover, the queries in mMARCO are derived from English-speaking users’ queries, which may not be as relevant to Spanish-speaking users (see section 4.1 for a more detailed analysis).
One feature of our dataset is the sparsity of its judgments, with only about one relevance judgment per query, similar to the widely used MS MARCO dataset. This means that, as compared to datasets with more judgments per query, such as MIRACL, it is more likely that relevant passages are not identified as such, and that IR systems will retrieve more false negatives, that is, passages erroneously considered non-relevant.
Similarly, unlike MIRACL, our dataset does not include passages explicitly judged as negative (non-relevant). This is a limitation for the training of models that require negative examples, and implies that negative mining techniques would be needed to train models with MessIRve.
Nevertheless, the use of sparse, only-positive judgments is a common practice in IR research due to the challenges and high costs associated with obtaining more complete annotations. Previous research has shown that findings based on sparse judgments, like those of MS MARCO, align well with results from more comprehensive and costly judgments, suggesting that while more exhaustive annotations may change absolute scores, they are unlikely to affect qualitative conclusions regarding the ranking of IR systems (Zhang et al., 2021a).
The automated data collection approach was primarily driven by the need for a large-scale dataset that did not require the costly human annotations associated with datasets like MIRACL and MS MARCO. This approach sacrifices control over quality because it relies on both the reliability of the automatically extracted queries and the accuracy of Google Search featured snippets. While this method may be seen as less dependable than human annotations, we have conducted a quality assessment to ensure the dataset’s reliability, as described in section 3.3.
In the following section we perform a comparison between the datasets in terms of their topics.

### v2

@@ Abstract
Information retrieval (IR) is the task of finding relevant documents in response to a user query. Although Spanish is the second most spoken native language, there are few Spanish IR datasets, which limits the development of information access tools for Spanish speakers. We introduce MessIRve, a large-scale Spanish IR dataset with almost 700,000 queries from Google’s autocomplete API and relevant documents sourced from Wikipedia. MessIRve’s queries reflect diverse Spanish-speaking regions, unlike other datasets that are translated from English or do not consider dialectal variations. The large size of the dataset allows it to cover a wide variety of topics, unlike smaller datasets. We provide a comprehensive description of the dataset, comparisons with existing datasets, and baseline evaluations of prominent IR models. Our contributions aim to advance Spanish IR research and improve information access for Spanish speakers.
@@
Francisco Valentini1, Viviana Cotik2,1, Damián Furman1, Ivan Bercovich3, Edgar Altszyler4, Juan Manuel Pérez1 1CONICET-Universidad de Buenos Aires. Instituto de Ciencias de la Computación (ICC). Buenos Aires, Argentina 2Universidad de Buenos Aires. Facultad de Ciencias Exactas y Naturales. Departamento de Computación. Buenos Aires, Argentina. 3University of California, Santa Barbara. California, U.S.A. 4Quantit {fvalentini,jmperez,vcotik}@dc.uba.ar, damian.a.furman@gmail.com, ibercovich@ucsb.edu, edgar@bequantit.com
@@ 1 Introduction
Given a user query, Information Retrieval (IR) is the process of finding and ranking relevant documents from a collection of data. IR systems have become increasingly important as they are the backbone for the currently widespread Retrieval Augmented Generation (RAG) systems, in which IR provides relevant passages that are subsequently fed into a Large Language Model (LLM, lewis2020retrieval; izacard2023atlas; ram2023incontext; shi2024replug). This approach has been shown to improve the quality of generated text, reducing the likelihood of factual errors or hallucinations by LLMs in several tasks (cheng2023uprise; jiang2023active; cheng2024lift; lin2024radit).
Natural language processing relies heavily on evaluation datasets, which establish a common ground for comparing algorithms and guide the development of new methods. While a wealth of resources are available for English, other languages often rely on multilingual models and datasets.
When addressing tasks in Spanish, resources are notably scarce. Spanish is the the second most spoken language in the world by number of native speakers, with over 486 million speakers. It is also the official language of 20 countries and is widely spoken in the United States (eberhard2024ethnologue). Despite its prevalence, at the time of writing this paper, the retrieval section of the popular MTEB benchmark (muennighoff-etal-2023-mteb) includes datasets for English, Chinese, French, and Polish, but omits Spanish.111huggingface.co/spaces/mteb/leaderboard
The scarcity of evaluation datasets is a major challenge for the development of IR systems for Spanish. Existing Spanish IR datasets are limited and present shortcomings in terms of dataset size, reliance on translations, restricted topical coverage, or lack of dialectal diversity. Which systems are best for Spanish IR? This key question remains unanswered due to the lack of a comprehensive Spanish IR benchmark.
@@
    Introducing MessIRve, a new large-scale dataset in Spanish that addresses current datasets’ limitations through automated data collection, the use of natural, non-translated queries, and the inclusion of dialectal diversity across the Spanish-speaking countries. [^2] MessIRve means works for me in Spanish (“me sirve”). The reference to Messi, player of the most popular sport in Spanish-speaking countries, football, also stresses the importance of using topics that are relevant to Spanish speakers.
@@
    Providing baseline evaluations of relevant models on the new dataset and the existing ones.
@@
    Making the dataset and a fine-tuned baseline model publicly available, so that other researchers can use them to develop and evaluate their own IR systems for Spanish. [^3] Dataset: hf.co/datasets/spanish-ir/messirve
    Model: hf.co/spanish-ir/multilingual-e5-large-ft-messirve-full
    Code: github.com/ftvalentini/MessirveSpanishIR
These contributions aim to spur further research in IR for the Spanish language and to facilitate the development of efficient information access tools for Spanish speakers.
@@ 2 Related work
mMARCO
8 841 823 532 761 1.00 - - 6 979 1.07
SQAC
6 247 14 934 1.01 1 861 1.00 1 908 1.00
MIRACL
10 373 953 2 162 9.96 648 9.94 1 515 9.95
PRES
10 037 - - - - 167 7.72
Multi-EuP
(yang-etal-2023-multi-eup)
2 371 - - - - 633 3.75
MessIRve (full) 14 047 759 537 730 1.06 - - 156 528 1.02
@@
In this section, we review well-documented, publicly available Spanish IR datasets. Table 1 presents their statistics. A dataset consists of queries, documents, and relevance judgments (thakur2021beir). Queries express information needs, documents contain potential answers, and relevance judgments assess document-query relevance. The full document set is the corpus.
The multilingual MSMARCO (mMARCO, bonifacio2021mmarco) translates the well-known MSMARCO dataset (bajaj2016ms) into Spanish using machine translation. Though not originally in Spanish, it provides a large-scale resource for training IR models, leveraging Bing search queries. Annotators judged the relevance of 10 Bing-retrieved passages per query, which form the document collection.
The Spanish Question Answering Corpus (SQAC, gutierrez2022maria) is a non-translated dataset inspired by SQUAD (rajpurkar-etal-2016-squad). Spanish-speaking annotators generated up to five questions per passage from Spanish Wikipedia and the AnCora corpus (taule2008ancora), encouraging synonym use. While lacking a full document collection, the set of relevant passages can serve as an IR corpus.
MIRACL (Multilingual Information Retrieval Across a Continuum of Languages, zhang2023miracl) includes Spanish and 17 other languages, emphasizing high-quality native speaker relevance judgments. Some annotators created queries from Wikipedia paragraphs, and others assessed the relevance of 10 passages retrieved by different models. Despite containing relatively few queries, MIRACL benefits from having more annotations per query and also negative relevance judgments.
Beyond multi-domain datasets, two domain-specific ones exist. The Spanish Passage Retrieval dataset (PRES, kamateri2019test) has health-related queries and articles from reliable Spanish health websites as documents. Relevance judgments were done via pooling, with annotators reviewing top-ranked results by different IR systems.
Finally, the Multilingual European Parliament Dataset (Multi-EuP, yang-etal-2023-multi-eup) uses European Parliament debate topics as queries, with speeches by Spanish parliamentarians as documents. Relevance judgments indicate whether a speech was part of a debate on a given topic.
@@ 3.1 Data collection
MessIRve is built by using questions from Google’s autocomplete API [^4] Endpoint: www.google.com/complete/search? as queries, and answers from Google Search “featured snippets” that link to Wikipedia as relevant documents. This data collection strategy is inspired by the GooAQ dataset (khashabi-etal-2021-gooaq-open).
@@
Starting with 120 predefined prefixes such as “qué” (what), “cómo” (how), “dónde” (where), “quién” (who), etc., we used the Google autocomplete API to obtain popular queries in Spanish starting with these prefixes (refer to Appendix C, Table 3 for the list of prefixes). We then iteratively expanded the set of prefixes based on the returned results until we approximately reached a predefined number of results. The API does not always return queries that begin strictly with the provided prefixes, resulting in a dataset where not all queries begin with the initial prefixes (see further details on query extraction in Appendix C).
@@
To find relevant passages for queries, we obtained Google Search “featured snippets” sourced from Wikipedia. These snippets are text excerpts displayed at the top of the search results in response to a query, deemed by Google to be relevant to the user’s information need.555blog.google/products/search/reintroduction-googles-featured-snippets
@@
We built a Spanish Wikipedia corpus with each paragraph corresponding to a single document. For this we used WikiExtractor (attardi2015wikiextractor) to process the Spanish Wikipedia dump of 2024-04-01 and matched the extracted Wikipedia passages to the valid query-document pairs obtained from Google Search featured snippets.
@@ 3.2 Data splits
Country Code #Unique Q #Unique Q #Unique Rel. Docs #Unique Rel. Articles Avg Q Length Avg Rel. Doc Length
- full 537 730 156 528 63 557 47 127 5.7 79.9
- none 356 040 101 359 44 869 34 306 5.8 80.8
Argentina ar 22 560 5 481 3 829 3 498 5.4 80.4
Bolivia bo 24 912 4 810 3 230 2 866 5.3 79.7
Chile cl 22 486 5 408 3 694 3 381 5.4 79.3
Colombia co 25 914 5 667 3 845 3 464 5.6 79.8
Costa Rica cr 23 662 5 690 4 047 3 693 5.5 79.3
Cuba cu 22 071 4 787 3 374 3 071 5.4 80.9
Dominican Rep. do 27 830 5 359 3 725 3 320 5.6 79.8
Ecuador ec 27 599 6 074 4 214 3 734 5.9 81.1
Spain es 23 476 7 148 5 004 4 654 5.5 80.0
Guatemala gt 22 971 4 630 3 091 2 755 5.5 79.5
Honduras hn 26 818 5 608 3 890 3 540 5.5 78.8
Mexico mx 32 258 8 099 5 563 4 991 5.5 79.2
Nicaragua ni 28 179 5 787 3 926 3 539 5.4 78.8
Panama pa 25 784 5 777 4 082 3 700 5.5 79.8
Peru pe 25 877 5 458 3 784 3 443 5.5 79.6
Puerto Rico pr 26 609 6 343 4 347 3 930 5.4 81.2
Paraguay py 24 885 5 306 3 785 3 435 5.4 78.7
El Salvador sv 25 935 5 806 3 999 3 619 5.5 78.3
United States us 23 498 4 234 2 818 2 572 5.6 81.9
Uruguay uy 20 902 5 525 4 062 3 701 5.3 76.9
Venezuela ve 27 123 5 733 3 924 3 455 5.5 81.5
Table 2: The MessIRve dataset. #: number of, Q: queries, Rel.: relevant, Doc.: document. Queries come from Google’s autocomplete API and documents are Wikipedia paragraphs. One paragraph may be relevant to multiple queries. The full set includes queries not specific to any country (code=none) plus the set of unique queries from all countries. Each query has one relevant document in each country. In the full set, a query may have multiple relevant documents. Statistics for the training set are nearly identical to those of the test set and are thus omitted.
The dataset is split into training and test queries in such a way that the Wikipedia article to which any relevant paragraph belongs is present in only one of the splits. The partitioning was done by country, with about 20% of the articles assigned to the test set. Splitting by article aims to reduce the overlap in topics between the training and test sets. We do not provide a validation or development set to allow users flexibility in their validation strategies, such as using cross-validation instead of relying on a single validation set.
Queries not specific to any country (label none) were combined with the set of unique queries from all countries to form a full set, which is also split into training and test sets. Unlike the country-specific sets, in the full set some queries can have multiple relevant documents because the same query may return different featured snippets in different country domains. Statistical information about the dataset splits is provided in Table 2. Representative query-document examples for each country are provided in Appendix C, Table 6.
@@ 3.3 Quality assessment
We evaluated the quality of the dataset using a method inspired by rybak-2023-maupqa, sampling query-document pairs from the full test set and manually rating them on three binary criteria:
@@
    Document relevance: Does the document help answer the query? Even if the passage does not provide a thorough answer, it can be deemed relevant if it contains information that, when combined with other sources, could help answer the query.
Three independent native Spanish speakers hired via Prolific (prolific.com) rated the same 100 random samples for 13.60 USD/hr. Considering the majority vote, 93% of the queries were rated as correct, 88% as unambiguous, and 97% of the documents were considered relevant to their queries. These results show that the dataset’s quality is acceptable, as the queries are generally clear and the documents annotated as relevant are most likely to contain information that helps answer the query.
Because these three criteria are intrinsically ambiguous, i.e. they depend on each rater’s interpretation, we measured the inter-rater agreement as the % of the queries on which all annotators agree. We obtained 80% for correctness, 65% for unambiguity, and 83% for relevance. These results support the reliability of the quality assessment.
To further assess the quality of the dataset, we used GPT-4o and Claude-3.5-Sonnet to annotate the relevance of other 650 query-document pairs. They identified 86% and 93.4% of the documents as relevant, respectively, with an agreement in 91.1% of the samples. To better understand the results, we manually reviewed the cases where both models marked documents as non-relevant. No clear pattern emerged from this analysis. Refer to Appendix B for examples.
Finally, as an additional validation, four of the paper’s authors carried out a quality assessment of these 650 pairs. We obtained 93.5% for correctness, 90.1% for unambiguity, and 90.2% for relevance. The consistence across independent annotators, coauthors, and LLM judges highlights the robustness of the dataset’s quality. Further details and a discussion of inter-rater agreement can be found in Appendix B.
@@ 4 Comparison with existing datasets
MessIRve provides detailed documentation of its data collection process, unlike existing open-domain datasets like MIRACL and SQAC, which lack specifics on Wikipedia passage selection. Our dataset is also the only one to account for Spanish dialectal diversity, whereas other datasets either ignore this or fail to document it. To validate this, we identified distinctive words in each country’s queries using log odds ratios (Table 4 in Appendix C). Examples for the countries with most speakers include “green” (USA, Green Card), “checar” (Mexico, “to check”), “frailejones” (Colombia, plant), “euskera” (Spain, Basque language), and “CUIT” (Argentina, tax ID).
@@
MessIRve is also much larger than other datasets (Table 1). While mMARCO is large, it relies on translations of English-speaking users’ queries, introducing translation errors and contents less relevant to Spanish-speaking users (see Section 4.2).
Like MSMARCO, MessIRve has sparse judgments, with only about one per query. It is thus more likely to have more false negatives compared to denser datasets like MIRACL. Additionally, MessIRve lacks explicitly negative judgments, requiring negative mining techniques for training (Section 5). IR systems trained using our dataset would benefit from more complete annotations.
Despite this, sparse and positive-only judgments are common in IR due to the challenges and high costs of annotation (lin2022pretrained). Explicit negative annotations are rare; for example, only one dataset in BEIR has them (Touché-2020) and they are not included in the benchmark (thakur2021beir). In addition, the number of negative annotations per query is usually small, e.g. MIRACL has 5.3 negative annotations per query, so some negative sampling is still required for training. Previous research shows that sparse judgments, like MSMARCO’s, yield rankings comparable to those from more comprehensive and costly annotations (zhang2021mr).
MessIRve remains valuable, especially within a broader evaluation framework combining datasets with varying judgment densities, as seen in BEIR (thakur2021beir). Our fine-tuning experiments (Section 5) also show that BM25-mined hard negatives improve performance; more advanced sampling strategies could further improve results.
Our automated data collection prioritized scale and avoiding costly manual annotations, like in MIRACL and MSMARCO. This sacrifices some control over quality, relying on automated query extraction and the accuracy of Google Search snippets. For this reason, we assessed the dataset’s reliability in Section 3.3. In the following sections we compare the datasets in terms of their format and topics.
4.1 Query format
Figure 1: Distribution of query lengths in each dataset. Boxplots exclude outliers for better visualization. Red crosses indicate the mean of each distribution.
MessIRve queries are generally shorter and less variable in length than those in other datasets (Figure 1). The shorter length of MessIRve queries is likely a characteristic of search engine queries, in contrast to datasets such as MIRACL or SQAC, which are collected from human annotators.
Short queries are not necessarily simple or generic. To show this, we computed the % of queries starting with typical question words in open-domain datasets (Spanish for “who”, “what”, “which”, “when”, “where”, “how many”, “how much”, “why”, “how”). We obtained 65.5% for MessIRve, 92% for MIRACL, 82.2% for SQAC, and 59.7% for mMARCO. This suggests that MessIRve is not disproportionately biased toward simple or generic questions when compared to other open-domain IR datasets. Examples of MessIRve queries not starting with these question words include “para qué sirve” (what is … for), “para qué es” (what is … for), and “en qué consiste …” (what does … consist of). These account for 2.33% of MessIRve queries but less than 0.30% in other datasets.
To further examine query structure, we measured prefix diversity using the entropy for prefix lengths from 1 to 6 (Appendix C, Figure 5). MessIRve and mMARCO show the highest entropy for multi-word prefixes, indicating greater variability in question openings. For single-word prefixes, MEUP ranks highest, which is consistent with its queries being debate topics.